### v1

# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

Abstract
This paper develops a unified framework of high-order accumulative regularization (AR) framework for convex and uniformly convex gradient norm minimization. Existing high-order methods often exhibit a gap: the function value residual decreases fast, while the gradient norm converges much slower. To close this gap, we introduce AR that systematically transforms fast function value residual convergence rate into fast (matching) gradient norm convergence rate.
Specifically, for composite convex problems, for computing an approximate solution such that the norm of its (sub)gradient does not exceed $\varepsilon,$ the proposed AR methods match the best corresponding convergence rate for the function value residual. We further extend the framework to uniformly convex settings, establishing linear, superlinear and sublinear convergence of the gradient norm under different lower curvature conditions. Moreover, we design parameter-free algorithms that require no input of problem parameters, e.g., Lipschitz constant of the $p$-th order gradient, the initial optimality gap and the uniform convexity parameter, and allows inexact solution for each high-order step. To our best knowledge, no parameter-free methods can attain such a fast gradient norm convergence rate which matches that of the function value residual in the convex case, and no such parameter-free methods for uniformly convex problems exist. These results substantially generalize existing parameter-free and inexact high-order methods and recover first-order algorithms as special cases, providing a unified approach for fast gradient minimization across a broad range of smoothness and curvature regimes.
keywords:
Smooth optimization, high-order methods, gradient norm minimization, Hölder condition, convex optimization, uniform convexity, parameter-free
1 Introduction
In this paper, we first consider the following unconstrained convex optimization problem:
(1) $\min_{x\in\mathbb{R}^{n}}f(x),$
where $f$ is proper and closed. We assume that $f(\cdot)$ is $p$-times differentiable with a Lipschitz continuous $p$-th derivative:
(2) $\|D^{p}f(x)-D^{p}f(y)\|\leq L_{p+1}(f)\|x-y\|,\quad\forall x,y\in\mathbb{R}^{n},$
where $\|D^{p}f(x)-D^{p}f(y)\|=\max\limits_{h\in\mathbb{R}^{n}}\{|D^{p}f(x)[h]^{p}-D^{p}f(y)[h]^{p}|:\,\|h\|\leq 1\}.$ We then extend the analysis to composite and uniformly convex problems.
The last decade has witnessed an increasing interest in the design and analysis of high-order methods for convex programming. High-order methods exploit high-order derivative information to achieve faster convergence compared to first-order methods. Nesterov and Polyak established the first global worst-case complexity analysis for the second-order schemes by designing a cubic regularization of Newton method (CNM) [30]. They showed that CNM takes at most $\mathcal{O}(\varepsilon^{-1/2})$ iterations to reduce the functional residual below a given precision $\varepsilon,$ i.e., $f(x)-f^{*}\leq\varepsilon,$ where $f$ is a twice-differentiable convex function with a Lipschitz continuous Hessian. Later, Nesterov obtained an improved complexity bound of $\mathcal{O}(\varepsilon^{-1/3})$ by developing an accelerated version of CNM [23]. After that, Monteiro and Svaiter [22] derived an improved iteration complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1})$ through an accelerated Newton proximal extragradient (ANPE). Later, a lower bound of $\Omega(\varepsilon^{-2/7})$ for second-order methods was established [3], which implies ANPE [22] is optimal up to a logarithmic factor. When it comes to $p>2$, the $p$-th order tensor method and its accelerated version are developed for convex functions with Lipschitz continuous $p$-th derivatives (2) in [5, 25]. Specifically, to achieve $f(x)-f^{*}\leq\varepsilon,$ the $p$-th order tensor method takes up to $\mathcal{O}(\varepsilon^{-1/p})$ iterations, while its accelerated version only requires $\mathcal{O}(\varepsilon^{-1/(p+1)})$ iterations [25]. Subsequent methods improved the iteration complexity to $\mathcal{O}(\varepsilon^{-2/(3p+1)}\log\varepsilon^{-1})$, such as those built upon the ANPE method [19, 6, 12], and the inexact high-order proximal-point methods [26, 28]. Recent methods [7, 20] managed to remove the extra logarithmic factor and thus are optimal.
On the other hand, the worst-case function for the class of functions with Lipschitz continuous Hessian has a discontinuous third derivative [28]. Therefore, quite interestingly, some algorithms that rely solely on second-order information can surpass the classical lower bound of $\Omega(\varepsilon^{-2/7})$ for the smaller function class whose second and third derivatives are Lipschitz continuous. For example, a second-order method only requires an iteration complexity of $\mathcal{O}(\varepsilon^{-1/4}),$ at each iteration, it calls the second-order oracle once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most, see Nesterov [27]. Furthermore, in [28], Nesterov shows that a second-order implementation of the third-order accelerated proximal-point method with segment search can achieve an iteration complexity of $\mathcal{O}(\varepsilon^{-1/5}),$ at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most. Following this, second-order methods have achieved the convergence rate of third-order accelerated tensor methods in terms of function value residual for composite problems [2, 1].
Furthermore, the aforementioned methods assume the access to the Lipschitz constant of the corresponding derivative, which is difficult to estimate in practice, and most assume exact solutions for each subproblem step, which is hard to obtain when $p>3$ [25]. Therefore, it is important to develop parameter-free and inexact algorithms that automatically estimate such parameters and solve each step inexactly, while maintaining similar theoretical guarantees. Among the second-order methods, Cartis et al. proposed an adaptive cubic regularized Newton method [8], and further relaxed the criterion for solving each subproblem while maintaining the convergence properties, which is widely used due to its numerical efficiency. Recent works [14, 15, 18, 13, 16] develop accelerated high-order methods that are fully independent of problem constants while maintaining superior theoretical iteration bounds, they achieve an iteration complexity of $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{1}{{p+1}}}}}),$ where $L_{0}$ is the initial line search value and $\theta$ represents the inexactness level of each subproblem.
Despite these progresses of high-order methods for convex optimization, the convergence guarantees in existing works are mainly established for function value residual, i.e., generating a point $\widehat{x}$ such that $f(\widehat{x})-f^{\ast}\leq\varepsilon$, where $f$ is the objective function, $f^{\ast}$ is its optimal value, and $\varepsilon>0$ is a given precision; not on the gradient norm criterion, i.e., $\|\nabla f(\widehat{x})\|\leq\varepsilon.$ While the former is theoretically appealing, it is difficult to verify in practice since $f^{*}$ is unknown. In contrast, the gradient norm criterion is easily computable, may serve as a measure of optimality. Moreover, the gradient norm minimization perspective also provides a practical mechanism to monitor the progress of the algorithm and can be used to estimate problem parameters adaptively. There are some works provided complexity results for the gradient norm criterion; however, they do not match the guarantees for function value residual. For example, in [14], the authors studied a regularized accelerated tensor method and established a complexity of $\mathcal{O}(\varepsilon^{-1/(p+1)}\log\varepsilon^{-1})$, which is worse than the corresponding function value residual complexity of the accelerated tensor methods [25] by a logarithmic factor. As another example, the ANPE method [22], exhibits a gradient norm complexity of $\mathcal{O}(\varepsilon^{-1/3}\log\varepsilon^{-1})$, which is worse than its nearly optimal function value residual complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1}).$ A third example is parameter-free and inexact method, where the best known iteration complexity for gradient norm minimization is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{p+1}{{p(p+2)}}}}})$ [15, 16]. However, this complexity is worse than the function value residual complexity, which is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}\varepsilon^{-\frac{1}{p+1}}).$
To summarize, the convergence rates in terms of gradient norm for different types of high-order methods do not match with the corresponding function value residual. This raises a natural question: is there a unified approach to translate the fast function value residual rate into a matching rate for gradient norm minimization?
To address this question, we propose a unified accumulative regularization (AR) method for gradient minimization and use it to accelerate existing higher-order algorithms, yielding best known gradient norm convergence rates for convex problems. Specifically, for an existing high-order algorithm $\mathcal{A}$, we utilize its fast function value residual decay together with its slow gradient norm decay, and design new gradient norm minimization algorithms that match the convergence rate of the function value residual. It is worth noting that our approach is inspired by the first-order gradient minimization algorithm of Lan et al. [21], and recovers their results when $p=1$. Our contributions can be briefly summarized as follows.
First, for a twice differentiable convex function with a Lipschitz continuous Hessian, we refine the estimating sequence analysis of the accelerated CNM (ACNM) to obtain a slow gradient norm convergence rate of $\mathcal{O}(\varepsilon^{-1/2})$, together with its original fast function residual decrease $\mathcal{O}(\varepsilon^{-1/3})$. Then we design a third-order AR method that uses ACNM as a subroutine. The method restarts intermittently, with each epoch initialized with the output from the previous epoch. Moreover, it applies a single accumulative regularization scheme whose parameter depends on $\varepsilon$ at initialization and then increases geometrically across epochs. We show that, to achieve $\|\nabla f(\widehat{x})\|\leq\varepsilon$, the number of iterations reduces to $\mathcal{O}(\varepsilon^{-1/3})$, thereby matching the complexity for the function value residual.
Second, we generalize the method to solve the composite convex problem
(3) $\min\limits_{x\in\mathbb{R}^{n}}\left\{f(x):=g(x)+h(x)\right\},$
where $g$ and $h$ are proper closed convex functions, $h$ maybe nonsmooth, and $g$ is $p$-times continuously differentiable on $\mathbb{R}^{n}.$ We assume that there exists at least one optimal solution $x^{*}.$ Furthermore, the level of smoothness of $g$ is characterized by the family of Hölder constants, i.e., $\exists\,p\geq 1,\ \nu\in[0\,\,1],\ L_{p,\nu}(g)>0$ such that:
(4) $\|D^{p}g(x)-D^{p}g(y)\|\leq L_{p,\nu}(g)\|x-y\|^{\nu},\quad\forall x,y\in\mathbb{R}^{n}.$
[…]
  1. (a)
    Suppose $g$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2,$ by choosing $\mathcal{A}$ as the $p$-th order accelerated tensor method [16], the $(p+\nu)$-th order AR framework only requires $\mathcal{O}(\varepsilon^{-1/{(p+\nu)}})$ iterations, which matches that of the function residual [25].
  2. (b)
    Suppose $g$ satisfies (4) with general $p\geq 1,$ and $\nu=1$, by choosing $\mathcal{A}$ as the nearly optimal tensor methods [19, 6, 12] or $p$-th order inexact proximal point method [28, 1], the $(p+1)$-th order AR framework requires $\mathcal{O}\left(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon}\right)$ iterations, which matches that of the function value residual. The extra logarithmic term disappears if $\mathcal{A}$ is chosen as the optimal $p$-th order tensor method (for function-residual) [7, 20] and can output a slow-gradient norm.
  3. (c)
    Suppose $g$ satisfies (4) with $p=3,\nu=1,$ by choosing $\mathcal{A}$ as the second-order method [28], the fourth-order AR framework achieves an iteration complexity of $\mathcal{O}(\varepsilon^{-{1}/{5}}),$ at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
Third, we develop inexact and parameter-free high-order AR framework for unconstrained convex optimization (1) where $f$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2.$ By choosing $\mathcal{A}$ as adaptive and inexact tensor methods [13, 16], the $(p+\nu)$-th order inexact and parameter-free AR framework only requires
$\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}\text{dist}(x_{0},X^{*})^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
calls to the $p$-th order oracle, where $L_{0}$ is the initial line search value and $\theta$ represents the inexactness level of the subproblem of each iteration. It does not require the knowledge of the number of iterations to run in advance, the number of restarting epochs, the knowledge of $L_{p,\nu}(f),$ and the initial optimality gap $\text{dist}(x_{0},X^{*})$. It matches the best known convergence rate in terms of function value residual for adaptive and inexact methods [18, 13, 16].
Fourth, we consider the composite convex problem (3), additionally assuming that $f$ is uniformly convex of order $q$ with parameter $\sigma_{q}:$
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\tfrac{\sigma_{q}}{q}\|y-x\|^{q},\quad\forall\,\,\,x,y\in\mathbb{R}^{n}.$
[…]
$\mathcal{O}\left(\left(\tfrac{(p+1)L_{p+1}(g)}{\sigma_{p+1}(f)}\right)^{\frac{2}{3p+1}}\log\tfrac{\|\nu_{0}\|}{\varepsilon}\right)$
iterations if $q=p+1,$ where $\nu_{0}\in\partial f(x_{0}).$ Notice that when $p=1$, this recovers the gradient minimization results for strongly convex and smooth function [21]. Furthermore, it improves the condition number dependence established in [10, 9] for linear decay of the function value residual. The linear rate improves to super linear rate if $q<p+1,$ i.e., $\|\widehat{\nu}\|\leq\varepsilon$ within at most
$\mathcal{O}\left({\log\log\left[\tfrac{\sigma_{q}(f)}{{q\varepsilon}}{\left(\tfrac{\sigma_{q}(f)}{qL_{p+1}(g)}\right)^{\frac{q-1}{p+1-q}}}\right]}+\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}\right)$
iterations. When $p=q=2,$ the algorithm achieves the iterations complexity of $\mathcal{O}({\log\log{\tfrac{\sigma^{2}_{2}(f)}{L_{3}(g)\varepsilon}}}+[{L_{3}(g)\|\nu_{0}\|}/{\sigma^{2}_{2}(f)}]^{{\frac{2}{7}}}),$ which nearly matches the lower bound [3] for function value residual. The algorithm become sublinear if $q>p+1,$ and requires
$\mathcal{O}\left(\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q}{\sigma_{q}(f)}\right)^{{\frac{2(q-1-p)}{(3p+1)(q-1)}}}{\varepsilon}^{-{\frac{2(q-1-p)}{(3p+1)(q-1)}}}\right)$
iterations. It improves over the sublinear rate $\mathcal{O}(\varepsilon^{-\frac{q-p-1}{pq}})$ when $q>p+1$ for function value residual [9].
Lastly, we develop inexact and parameter-free gradient minimization which do not require the knowledge of Lipschitz continuous parameter or uniform convex parameter. Using unconstrained uniformly convex optimization problem with $q=p+1$ as a showcase, our parameter-free and inexact method requires
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right)$
iterations to compute a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$. It is the first parameter-free and inexact method that does not require the knowledge of the uniform convexity parameter $\sigma_{p+1}(f),$ while still matching the fastest known convergence rate of methods that assume $\sigma_{p+1}(f)$ is known. A very recent work [34] derive similar complexity as the prior works [10, 9] and established global and local complexities with linear ($q=p+1$) and superlinear ($q>p+1$) decay by an inexact and adaptive approach. Notice that it still assume the access to the uniform convexity parameter.
To the best of our knowledge, all of the above gradient-norm convergence results are new. The only exception is the recent work [11], which established a complexity of $\mathcal{O}(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon})$ for convex problem. However, the result appears to be more restrictive as they require $p\geq 2,$ thus hard to expand to the cases of bounded Hessian $(p=2,\nu=0)$ or $L$-smooth $(p=1,\nu=1).$ Moreover, it is restricted to the unconstrained case and does not extend to general convex composite or uniformly convex objectives. Unlike AR, it relies on known problem parameters and requires the subproblem of each iteration to be solved exactly.
The rest of this paper is organized as follows. In Section 2, we provide a brief review of the ACNM method and its modified analysis for convex functions with smooth Hessians. Then, we construct a third-order AR approach to accelerate the gradient norm convergence rate. In Section 3, we propose a general AR framework and employ different types of subroutines $\mathcal{A}$ for structured convex composite problems. In Section 4, we derive parameter-free inexact high-order AR framework for convex functions. In Section 5, we further investigate gradient norm minimization for uniformly convex functions and develop a parameter-free and inexact algorithm.
1.1 Notation and terminology
We use $\|\cdot\|$ to denote the Euclidean norm in $\mathbb{R}^{n},$ which is associated with the inner product $\langle\cdot,\cdot\rangle.$ Denote the directional derivative of function $f$ at $x$ along directions $h_{i}\in\mathbb{R}^{n},i=1,\dots,p$ as $D^{p}f(x)[h_{1},\dots,h_{p}].$ Note that $D^{p}f(x)[\cdot]$ is a symmetric $p$-linear form. For example, for any $x\in\text{dom }f$ and $h_{1},h_{2}\in\mathbb{R}^{n},$ we have
$Df(x)[h_{1}]=\langle\nabla f(x),h_{1}\rangle\quad\text{and}\quad D^{2}f(x)[h_{1},h_{2}]=\langle\nabla^{2}f(x)h_{1},h_{2}\rangle.$
Its norm is induced by the Euclidean norm.
$\|D^{p}f(x)\|\coloneqq\max_{h_{1},\dots,h_{p}}\left\{D^{p}f(x)[h_{1},\dots,h_{p}]:\|h_{i}\|\leq 1,i=1,\dots,p\right\}.$
For convex function $f,$ $\partial f(x)$ denotes set of subdifferential at $x.$ For any real number $s,$ $\lceil s\rceil,\lfloor s\rfloor$ denote the nearest integer to $s$ from above and below. Let $m]\triangleq\{1,\dots,m\}$, with $m\in\mathbb{N}_{+}.$
2 Accelerated Cubic Regularization of the Newton Method
In this section, we consider the unconstrained convex optimization problem (1), where $f$ is a twice differentiable convex function on $\mathbb{R}^{n},$ and it satisfies (2) with $p=2,$ i.e., there exists $L_{3}(f)>0$ such that $\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq L_{3}(f)\|x-y\|,$ for all $x,y\in\mathbb{R}^{n}.$
We start with a modified version of the ACNM [23], and show that unlike the ACNM which only outputs function value, the modified ACNM can output a slow gradient norm and the its original fast function value residual simultaneously. Define $T_{M}(x)\coloneqq\arg\min_{y\in\mathbb{R}^{n}}\xi_{2,x}(y)$, where
$\xi_{2,x}(y)={f(x)+\langle\nabla f(x),y-x\rangle+\tfrac{1}{2}\langle\nabla^{2}f(x)(y-x),y-x\rangle}+\tfrac{M}{6}\|y-x\|^{3},$
and $\arg\min$ indicates that $T_{M}(x)$ is chosen from the set of global minimizers of the corresponding problem.
Algorithm 1 Accelerated Cubic Regularization of Newton’s method
1: $x_{0}\in\mathbb{R}^{n},\{a_{k}\}_{k\geq 1},$ $a_{k}>0,A_{1}>0.$ Compute $x_{1}=T_{L_{3}}(x_{0})$ and define $f_{1}(x)\coloneqq{f(x_{1})+{\tfrac{1}{\sqrt{L_{3}(f)+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}}+\tfrac{C}{6}\|x-x_{0}\|^{3}.$
2: for $k\geq 1$ do
3:   Compute $\nu_{k}=\operatorname*{arg\,min\,}_{x\in\mathbb{R}^{n}}f_{k}(x).$ Choose $A_{k+1}=A_{k}+a_{k}>0$ and $y_{k}$ $=(1-\alpha_{k})x_{k}+\alpha_{k}\nu_{k},\quad\text{where}\quad\alpha_{k}=\tfrac{a_{k}}{A_{k+1}},\quad a_{k}>0.$
4:   Compute $x_{k+1}=T_{M}(y_{k})$ and update $f_{k+1}(x)$ $=f_{k}(x)+a_{k}\left(f(x_{k+1})+\langle\nabla f(x_{k+1}),x-x_{k+1}\rangle\right).$
5: $x_{k+1}.$
With a modified ACNM analysis, we can establish the following relations:
$\mathcal{R}_{k}^{1}:\quad$ $A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}\leq f_{k}^{*}\equiv\min\limits_{x\in\mathbb{R}^{n}}f_{k}(x).$
$\mathcal{R}_{k}^{2}:\quad$ $f_{k}(x)\leq A_{k}f(x)+\tfrac{(2L+C)\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{6}\|x-x_{0}\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}.$
Here $\mathcal{R}_{k}^{1}$ shows an lower bound for the minimum function value of the estimating function $f_{k}(x),$ and $\mathcal{R}_{k}^{2}$ establishes an upper bound for $f_{k}(x).$ See Appendix A for the proof. Thus, we have the following convergence guarantee for the modified ACNM.
Lemma 2.1.
Let the sequence $\{x_{k}\}^{\infty}_{k=1}$ be generated by Algorithm 1 with the parameters $M=2L_{3}(f),C={12L_{3}(f)}/{(\sqrt{2}-1)^{2}},a_{k}={(k+1)(k+2)}/{2},A_{1}=1,$ then for any $k\geq 1,$ we have
$f(x_{k})-f(x^{*})+\tfrac{1}{\sqrt{3L_{3}(f)}}\|\nabla f(x_{k})\|^{\frac{3}{2}}\leq\tfrac{80L_{3}(f)\|x_{0}-x^{*}\|^{3}}{k(k+1)(k+2)},$
where $x^{*}$ is an optimal solution to the problem (1).
It should be noted that the convergence rate of the gradient norm is slower than that of the function value residual. To achieve a point $\tilde{x}$ such that $\|\nabla f(\tilde{x})\|\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/2}\|x_{0}-x^{*}\|/\varepsilon^{1/2})$ iterations, while to achieve the function value residual $f(\tilde{x})-f^{*}\leq\varepsilon,$ Algorithm 1 requires $\mathcal{O}(L_{3}(f)^{1/3}\|x_{0}-x^{*}\|/\varepsilon^{1/3})$ iterations. Note that using a regularization technique could improve the convergence rate to $\mathcal{O}([L_{3}(f)^{1/3}\|x_{0}-x^{*}\|^{2/3}/\varepsilon^{1/3}]\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon}))$ [24]. However, the rates differ by a logarithmic term.
We next outline a third-order AR method that utilizes Algorithm 1 to improve its gradient norm convergence rate to that of the function value residual. Notably, the design of Algorithm 2 is built upon accumulative regularization, where the search points $\{x_{s}\}_{s\in[S]}$ are generated by inexactly solving a third-order proximal mapping defined in (7). Algorithm 2 requires a sequence of regularization parameters $\{\sigma_{s}\}_{s\in[S]}$, and the previous search points $\{x_{s}\}_{s\in[S-1]}$ are accumulated within the regularization term. Unlike the fixed regularization [24], Algorithm 2 uses an accumulative, restart-based scheme: each epoch starts at $x_{s-1}$, and the regularizer adds a new term centered at $x_{s}$ while preserving all earlier terms (rather than centered at $x_{0}$). Such an accumulative regularization is close in spirit to the classic accelerated proximal point method [17].
Observe that the proximal term is cubic. We next develop several properties of this cubic term for use in the subsequent algorithmic analysis. Denote the cubic function as $d_{3}(x)=\|x-x_{0}\|^{3}/3$, for any $x_{0}\in\mathbb{R}^{n}.$ Note that it has a Lipschitz continuous Hessian [32] as follows:
(5) $\|D^{2}d_{3}(x)-D^{2}d_{3}(y)\|\leq 4\|x-y\|,\quad\forall\,\,x,y\in\mathbb{R}^{n}.$
Furthermore, it is a uniformly convex function of degree $3$ with parameter $1/2$ [29].
(6) $d_{3}(x)-d_{3}(y)-\langle\nabla d_{3}(y),x-y\rangle\geq\tfrac{1}{6}\|x-y\|^{3},\quad\forall\,\,x,y\in\mathbb{R}^{n}.$
Algorithm 2 A third-order AR method for gradient minimization
1: Total number of epochs $S,$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0$ and initial point $x_{0}\in\mathbb{R}^{n}.$
2: $x_{S}.$
3: for $s=1,\dots,S$ do
4:   Set initialization at epoch $s$ to $x_{s-1}.$
5:   Compute an approximate solution $x_{s}$ of the proximal subproblem by running Algorithm 1 with the initialization $x_{s-1}$ for $N_{s}$ iterations. (7) $x_{s}\approx$ $\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x-{x}_{i-1}\|^{3}}{3}\right\}.$
6: output $x_{S}$
We now present the convergence rate of the third-order AR Algorithm 2.
Proposition 2.2.
Suppose the parameters are set to
(8) $S=\left\lceil\log_{4}\tfrac{L_{3}(f)D^{2}}{\varepsilon}\right\rceil+1,\,\,\sigma_{s}=\tfrac{4^{s-2}\varepsilon}{D^{2}},\,\,N_{s}=\left\lceil 4\left(\tfrac{480[L_{3}(f)+4\sigma_{s}]}{{\sigma_{s}}}\right)^{\frac{1}{3}}\right\rceil,$
where $D$ is an upper bound on the distance to the set of optimal solutions, i.e., $D\geq\min\limits_{x^{*}\in X^{*}}\|x_{0}-x^{*}\|.$ Then Algorithm 2 can compute an approximate solution $x_{S}$ such that $\|\nabla f(x_{S})\|\leq\varepsilon$ after at most
$\left\lceil\tfrac{128L_{3}(f)^{\frac{1}{3}}D^{\frac{2}{3}}}{{\varepsilon^{\frac{1}{3}}}}+128S\right\rceil$
evaluations of the first-, and second-order information of $f$.
Proof 2.3.
By (7), we have $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2}.$ We start with bounding $\|\nabla f_{S}(x_{S})\|.$ By the property of the power prox-function in (5), we conclude $f_{S}(\cdot)$ has a Lipschitz continuous Hessian with parameter $L_{3}(f_{S})=L_{3}(f)+4\sigma_{S}.$ For all $s\in[S],$ denote $x_{s}^{*}$ as its exact solution, by Lemma 2.1, we have
(9) $\|\nabla f_{S}(x_{S})\|$ ${\leq}\,{\left[3L_{3}(f_{S})\right]^{\frac{1}{3}}\left(\tfrac{80L_{3}(f_{S})\|x_{S}^{*}-x_{S-1}\|^{3}}{N_{S}^{3}}\right)^{\frac{2}{3}}},$
We proceed with bounding $\|x_{S}^{*}-x_{S-1}\|.$ By optimality conditions of (7) at $x_{s-1}^{*}$ and $x_{s}^{*},$ there holds
$f(x_{s-1}^{*})+\textstyle\sum_{i=1}^{s-1}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x_{s-1}^{*}-{x}_{i-1}\|^{3}}{3}+\tfrac{(\sigma_{s}-\sigma_{s-1})\|x_{s}^{*}-{x}_{s-1}\|^{3}}{3}$
$\leq f(x_{s}^{*})+\textstyle\sum_{i=1}^{s-1}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x_{s}^{*}-{x}_{i-1}\|^{3}}{3}+\tfrac{(\sigma_{s}-\sigma_{s-1})\|x_{s}^{*}-{x}_{s-1}\|^{3}}{3}$
(10) $\leq f(x_{s-1}^{*})+\textstyle\sum_{i=1}^{s-1}\tfrac{(\sigma_{i}-\sigma_{i-1})\|x_{s-1}^{*}-{x}_{i-1}\|^{3}}{3}+\tfrac{(\sigma_{s}-\sigma_{s-1})\|x_{s-1}^{*}-{x}_{s-1}\|^{3}}{3}.$
Thus, we have $\|x_{s}^{*}-x_{s-1}\|\leq\|x_{s-1}^{*}-x_{s-1}\|,$ for all $s\in[S].$ Using this fact, we can derive the linear convergence of the solution error per epoch, i.e., $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}^{*}-x_{s-1}\|/4$ as follows.
$\|x_{s}-x_{s}^{*}\|^{3}$ $\overset{\text{(a)}}{\leq}\tfrac{6[f_{s}(x_{s})-f_{s}(x_{s}^{*})]}{\sigma_{s}}\overset{\text{(b)}}{\leq}\tfrac{6}{\sigma_{s}}\tfrac{80L_{3}(f_{s})}{N_{s}^{3}}\|x_{s}^{*}-x_{s-1}\|^{3}$
(11) $\overset{\text{(c)}}{\leq}\tfrac{6}{\sigma_{s}}\tfrac{80L_{3}(f_{s})}{N_{s}^{3}}\|x_{s-1}^{*}-x_{s-1}\|^{3}\overset{\text{(d)}}{\leq}\tfrac{1}{64}\|x_{s-1}^{*}-x_{s-1}\|^{3},$
where in (a), we used the convexity of $f,$ the uniform convexity of power function in (6), and $\nabla f_{s}(x_{s}^{*})=0,$ in (b) we used Lemma 2.1, in (c), we used the fact $\|x_{s}^{*}-x_{s-1}\|\leq\|x_{s-1}^{*}-x_{s-1}\|,$ and in (d), we substitute the choice of $N_{s}$ in (8). Thus, substituting (2.3) and (8) into (9), we have
$\|\nabla f_{S}(x_{S})\|^{3}$ $\leq{3L_{3}(f_{S})\left(\tfrac{80L_{3}(f_{S})\|x_{S-1}^{*}-x_{S-1}\|^{3}}{N_{S}^{3}}\right)^{2}}\leq 3(L_{3}(f)+\tfrac{4^{S-1}\varepsilon}{D^{2}})(\tfrac{\varepsilon D}{96\times 4^{2S}})^{2}\leq\tfrac{\varepsilon^{3}}{8}.$
It remains to bound $\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2}.$
$\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{s}-{x}_{i-1}\|^{2}$
$=\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\bigg(2\textstyle\sum_{\ell>k\geq i}^{S}\langle x_{k}-x_{k-1},x_{\ell}-x_{\ell-1}\rangle+\textstyle\sum_{k=i}^{S}\left\|x_{k}-x_{k-1}\right\|^{2}\bigg)$
$\leq\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\bigg(2\textstyle\sum_{\ell>k\geq i}^{S}\|x_{k}-x_{k-1}\|\|x_{\ell}-x_{\ell-1}\|+\textstyle\sum_{k=i}^{S}\left\|x_{k}-x_{k-1}\right\|^{2}\bigg).$
By using the fact $\|x_{k}^{*}-x_{k-1}\|\leq\|x_{k-1}^{*}-x_{k-1}\|,$ we have
$2\textstyle\sum_{\ell>k\geq i}^{S}\|x_{k}-x_{k-1}\|\|x_{\ell}-x_{\ell-1}\|\leq 8\textstyle\sum_{\ell>k\geq i}^{S}\|x_{k-1}^{*}-x_{k-1}\|\|x_{\ell-1}^{*}-x_{\ell-1}\|$
[…]
Similarly, we have $\sum_{k=i}^{S}\left\|x_{k}-x_{k-1}\right\|^{2}\leq 2\sum_{k=i}^{S}{D^{2}}/{4^{2k}}.$ Therefore, we have
$\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2}\leq\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\left(\textstyle\sum_{k=i}^{S}\tfrac{2D^{2}}{4^{2k}}+\textstyle\sum_{k=i}^{S}\tfrac{D^{2}}{4^{2k-2}}\tfrac{8}{3}\right)$
[…]
Combining it with $\|\nabla f(x_{S})\|\leq\|\nabla f_{S}(x_{S})\|+\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}-{x}_{i-1}\|^{2},$ we conclude $\|\nabla f(x_{S})\|\leq\varepsilon.$ The total evaluations of the first-, and second-order information of $f$ is bounded by
$\textstyle\sum_{s=1}^{S}N_{s}$ $\overset{\textnormal{(e)}}{\leq}\textstyle\sum_{s=1}^{S}\left\lceil\tfrac{4[L_{3}(f)^{\frac{1}{3}}+\left({4^{s-1}\varepsilon}/{D^{2}}\right)^{\frac{1}{3}}]480^{\frac{1}{3}}}{\left({4^{s-2}\varepsilon}/{D^{2}}\right)^{\frac{1}{3}}}\right\rceil\leq\left\lceil\tfrac{128L_{3}(f)^{\frac{1}{3}}D^{\frac{2}{3}}+128S\varepsilon^{\frac{1}{3}}}{\varepsilon^{\frac{1}{3}}}\right\rceil,$
where in (e), we substitute the choices of $\sigma_{k},N_{k}$ in (8) and used the relation $(a+b)^{1/3}\leq a^{1/3}+b^{1/3},$ for all $a,b\geq 0.$ This concludes the proof.
The following comments are in order.
(i) On the parameter choices: The regularization parameter $\sigma_{s}$ is exponentially increasing, while the epoch length $N_{s}$ is exponentially decreasing, scaling inversely with respect to $\sigma_{s}$, i.e., $N_{s}=\Theta([L_{3}(f)/\sigma_{s}]^{1/3})$. The dominant iterations belongs to the first epoch, which determines the final convergence rate, as the remaining $N_{s}$ are summable. Furthermore, the number of epochs $S$ is chosen such that $4\sigma_{S}$ reaches $L_{3}(f)$, after which the algorithm will terminate.
(ii) On the convergence rate: Observe that the convergence rate in terms of the gradient norm in Proposition 2.2 removes the extra $\log({L_{3}(f)\|x_{0}-x^{*}\|^{2}}/{\varepsilon})$ in [24], and matches the convergence rate of the function value residual for ACNM [23]. Furthermore, the regularization parameter $\sigma_{s}$ increases geometrically as the algorithm approaches the true solution $x^{*}$, which enables the algorithm to use shorter epoch lengths $N_{s}$, thus achieving a faster overall convergence rate.
3 Gradient Minimization Framework for Structured Convex Problems
In this section, we consider solving the convex composite problem (3). We first introduce a general high-order AR framework, and then in Section 3.1, we study its convergence results when the smooth function part has $\nu$-Hölder continuous $p$-th derivatives. In Section 3.2, we focus on the case when the smooth function part has Lipschitz continuous $p$-th order derivative (i.e., $\nu=1$).
Consider the high-order AR framework as in Algorithm 3.
Algorithm 3 High-order AR framework for gradient minimization
1: Total number of epochs $S,$ strictly increasing regularization parameters $\{\sigma_{s}\}_{s=0}^{S}$ with $\sigma_{0}=0$ and initial point $x_{0}\in\mathbb{R}^{n}.$
2: $x_{S}.$
3: for $s=1,\dots,S$ do
4:   Compute an approximate solution $x_{s}$ of the proximal subproblem (12) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$ where $p+\nu\geq 2,$ by running some subroutine $\mathcal{A}$ with the initialization $x_{s-1}$ for $N_{s}$ iterations.
5: output $x_{S}$
Observe that Algorithm 3 shares the acceleration spirit of Algorithm 2, and the power of the regularization term depends on the smoothness level of $g.$ The subroutine $\mathcal{A}$ needs to satisfy some properties for AR to converge, for example, $\mathcal{A}$ outputs a slow gradient norm decay and a fast function value residual decay, which we will specify later in the convergence analysis.
Notice that the regularization terms in the subproblem (12) build upon the power prox-function $d_{p+\nu}:\mathbb{R}^{n}\rightarrow\mathbb{R},$ defined as $d_{p+\nu}(x)\coloneqq 1/{(p+\nu)}\|x-x_{0}\|^{p+\nu},x_{0}\in\mathbb{R}^{n}.$ We review its uniform convexity and $\nu$-Hölder continuity of the $p$-th order derivatives as follows [23, 32]. Notice that we only require $p+\nu\geq 2,$ therefore, AR can cover the first order algorithm for $L$-smooth convex function, i.e., $p=\nu=1,$ and the bounded Hessian case, i.e., $p=2,\nu=0.$
Lemma 3.1.
For all $\nu\in[0\,\,1],{p+\nu\geq 2},$ $d_{p+\nu}(x)$ is a uniformly convex function of degree $p+\nu$ with parameter $2^{-(p+\nu-2)}.$
(13) $d_{p+\nu}(x)-d_{p+\nu}(y)-\langle\nabla d_{p+\nu}(y),x-y\rangle\geq\tfrac{1}{p+\nu}\left(\tfrac{1}{2}\right)^{p+\nu-2}\|x-y\|^{p+\nu}.$
Furthermore, its $p$-th order derivatives satisfies $\nu$- Hölder continuity.
(14) $\|D^{p}d_{p+\nu}(x)-D^{p}d_{p+\nu}(y)\|\leq\tfrac{2}{p+\nu}\textstyle\prod_{i=1}^{p}(i+\nu)\|x-y\|^{\nu}.$
Suppose Algorithm 3 terminates at epoch $S$ with a point $x_{S}^{k},k\geq N_{S},$ then, by the construction of $f_{S},$ and suppose $p+\nu\geq 2,$ we have
$\nu_{S}^{k}\coloneqq v_{S}^{k}-\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|{x}_{S}^{k}-x_{i-1}\|^{p+\nu-2}({x}_{S}^{k}-x_{i-1})\in\partial f({x}_{S}^{k}),$
where $v_{S}^{k}\in\partial f_{S}(x_{S}^{k}).$ Therefore, the subgradient norm $\|\nu_{S}^{k}\|$ can be decomposed into two parts as follows.
(15) $\|\nu_{S}^{k}\|$ $\leq\|v_{S}^{k}\|+{\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}^{k}-{x}_{i-1}\|^{p+\nu-1}}.$
We next show that the second term can be properly controlled by the regularization parameters, provided that the inner subroutine satisfies a linear convergence condition $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ which can be satisfied by our subroutines, as we will demonstrate in the next subsections.
Denote $D$ as an upper bound on the distance to the set of optimal solutions, i.e.,
(16) $\text{dist}(x_{0},X^{*})\coloneqq\min_{x^{*}\in X^{*}}\|x_{0}-x^{*}\|\leq D.$
It characterizes the initial optimality condition and will be used frequently in our convergence analysis. We have the following convergence guarantee for (15).
Lemma 3.2.
Suppose $p+\nu\geq 2,$ the subroutine can compute $x_{s},$ for all $s\in[S]$, such that $\|x_{s}-x_{s}^{*}\|\leq\|x_{s-1}-x_{s-1}^{*}\|/4,$ and suppose in the last epoch $S$, for all $k\geq N_{S}$, there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq\|x_{S-1}-x_{S-1}^{*}\|/4.$ Then, for all $k\geq N_{S},$ there holds
$\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}^{k}-x_{i-1}\|^{p+\nu-1}\leq\tfrac{2\sigma_{S}D^{p+\nu-1}}{4^{(p+\nu-1)(S-2)}}+\tfrac{(18D)^{p+\nu-1}}{2}\textstyle\sum_{i=1}^{S}\tfrac{\sigma_{i}-\sigma_{i-1}}{4^{i(p+\nu-1)}},$
where $D$ is defined in (16).
Proof 3.3.
Similar to the proof of Proposition 2.2, by using the optimality condition of (12), for all $s\geq 1,$ $s\in[S],$ there holds
(17) $\|x_{s-1}-x_{s}^{*}\|$ $\leq\|x_{s-1}-x_{s-1}^{*}\|.$
Furthermore, for the last epoch $S,$ if $p+\nu\geq 2,$ there holds
$\big\|\textstyle\sum_{k=i}^{S-1}(x_{k}-x_{k-1})\big\|^{2}\leq 2\textstyle\sum_{k=i}^{S-1}\left(\left\|x_{k}-x_{k}^{*}\right\|^{2}+\left\|x_{k-1}^{*}-x_{k-1}\right\|^{2}\right)$
(18) $\quad\quad\quad\quad+\textstyle\sum_{\ell>k\geq i}^{S-1}\left(\|x_{k}-x_{k}^{*}\|+\|x_{k-1}^{*}-x_{k-1}\|\right)\left(\|x_{\ell}-x_{\ell}^{*}\|+\|x_{\ell-1}^{*}-x_{\ell-1}\|\right).$
By the convexity of the function $x\rightarrow\|x\|^{p+\nu-1},$ for $p+\nu\geq 2,$ we have
$\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}^{k}-{x}_{i-1}\|^{p+\nu-1}$
$\leq 2^{p+\nu-2}\left(\sigma_{S}\left\|x_{S}^{k}-x_{S-1}\right\|^{p+\nu-1}+\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\big\|\textstyle\sum_{k=i}^{S-1}(x_{k}-x_{k-1})\big\|^{p+\nu-1}\right).$
Inserting $x_{S}^{*}$ into the first part, and using the condition $\|x_{S}^{k}-x_{S}^{*}\|\leq(1/4)\|x_{S-1}-x_{S-1}^{*}\|,$ for all $k\geq N_{S},$ we have
   $\,\,\,2^{p+\nu-2}\sigma_{S}\|x_{S}^{k}-x_{S-1}\|^{p+\nu-1}$
$\,\,\leq 4^{p+\nu-1}\sigma_{S}\|x_{S}^{k}-x_{S}^{*}\|^{p+\nu-1}+4^{p+\nu-1}\sigma_{S}\|x_{S}^{*}-x_{S-1}\|^{p+\nu-1}$
[…]
Using (18), and the condition $\|x_{s}-x_{s}^{*}\|\leq\,(1/4)\|x_{s-1}-x_{s-1}^{*}\|,$ we have
$\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\big\|\textstyle\sum_{k=i}^{S-1}(x_{k}-x_{k-1})\big\|^{p+\nu-1}$
$\leq\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\bigg[\textstyle\sum_{k=i}^{S-1}\tfrac{4D^{2}}{16^{k-1}}+\textstyle\sum_{\ell>k\geq i}^{S-1}\tfrac{8D^{2}}{4^{k+\ell-2}}\bigg]^{\frac{p+\nu-1}{2}}$
$\leq(9D)^{p+\nu-1}\textstyle\sum_{i=1}^{S}\tfrac{\sigma_{i}-\sigma_{i-1}}{4^{i(p+\nu-1)}},$
which concludes the proof.
In the following, we specify Assumptions needed to satisfy the linear convergence in terms of the distance to the optimal solution across epochs, i.e., $\|x_{s}-x_{s}^{*}\|\leq\,(1/4)\|x_{s-1}-x_{s-1}^{*}\|,$ and show the convergence rate of Algorithm 3.
3.1 Extension from accelerated cubic newton to high-order tensor method
In this subsection, we study AR with tensor methods as subroutines to handle convex functions with $\nu$-Hölder smooth continuous $p$-th derivatives. We make the following assumption regarding the subroutine $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$
Assumption 1.
The approximate solution $x_{s}$ exhibits the following performance guarantees: after $N_{s}$ iterations of the subroutine $\mathcal{A}_{s},$ there holds
(19) $f_{s}({x_{s}})-f_{s}(x_{s}^{*})$ $\leq\tfrac{{C}_{\mathcal{A}}L_{p,\nu}(g)\|x_{s}^{*}-{x_{s-1}}\|^{p+\nu}}{(N_{s}-1)^{p+\nu}},\quad\forall\,\,1\leq s\leq S.$
For the $S$-th epoch, there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),N_{S}\leq k\leq 2N_{S},$ such that
(20) $\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|v_{S}^{k}\|\leq\tfrac{{C}_{\mathcal{A}}L_{p,\nu}(g)\|x_{S}^{*}-x_{S-1}\|^{p+\nu-1}}{(N_{s}-1)^{p+\nu-1}},$
where ${C}_{\mathcal{A}}>1$ is a universal constant that depends on $\mathcal{A}_{s}.$
The above Assumption holds for a wide range of algorithms with different $(p,\nu).$ See the end of this subsection for applications. We next demonstrate that with 1 and properly chosen epoch lengths $\{N_{s}\}_{s\in[S]},$ we can achieve the linear convergence in terms of the distance to the optimal solution across epochs, i.e., $\|x_{s}-x_{s}^{*}\|\leq\,(1/4)\|x_{s-1}-x_{s-1}^{*}\|.$
Lemma 3.4.
Suppose $p+\nu\geq 2,$ and for all $s\in[S],$ the epoch length satisfies
(21) $N_{s}\geq\left\lceil 4\left(\tfrac{2^{{p+\nu-2}}(p+\nu){C}_{\mathcal{A}}L_{p,\nu}(g)}{{\sigma_{s}}}\right)^{\frac{1}{p+\nu}}\right\rceil+1,$
where $D$ is defined in (16). If the subroutine $\mathcal{A}$ for solving subproblems (12) satisfies 1, then, for all $s\in[S],$ Algorithm 3 can compute $x_{s}$ such that
(22) $\|x_{s}-x_{s}^{*}\|\leq\tfrac{\|x_{s-1}-x_{s-1}^{*}\|}{4}.$
Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
Proof 3.5.
By the uniform convexity of $f_{s}$ and the optimality condition $\langle v_{s},x_{s}-x_{s}^{*}\rangle\geq 0,$ where $v_{s}\in\partial f_{s}(x_{s}),$ we have
(23) $f_{s}(x_{s})-f_{s}(x_{s}^{*})$ $\geq\tfrac{\sigma_{s}}{p+\nu}\left(\tfrac{1}{2}\right)^{p+\nu-2}\|x_{s}-x_{s}^{*}\|^{p+\nu},\quad\forall s\in[S].$
Combining it with (19) in 1 and (17) in Lemma 3.2, we have
[…]
Therefore, by the epoch length choice in (21), we have
[…]
Similarly, notice that in the last epoch, for all $N_{S}\leq k\leq 2N_{S},$ there holds
[…]
Substituting the epoch length condition (21) concludes the proof.
Denote $p$-th order oracle at point $x$ as $\{f(x),\nabla f(x),\nabla^{2}f(x),\dots,\nabla^{p}f(x)\},p\geq 1.$ The following theorem states the main convergence results of the high-order AR framework with subroutine $\mathcal{A}$ applied for convex composite problem (3).
Theorem 3.6.
Suppose the Assumptions in Lemma 3.4, and
(25) $S\coloneqq\left\lceil\log_{2^{p+\nu-1}}\tfrac{{C}_{\mathcal{A}}4^{(p+\nu-2)}D^{{p+\nu-1}}L_{p,\nu}(g)}{\varepsilon}\right\rceil+1,\,\,\sigma_{s}\coloneqq\tfrac{2^{(p+\nu-1)(s-1)}\varepsilon}{{C}_{\mathcal{A}}4^{(p+\nu-2)}D^{p+\nu-1}},$
for all $s\in[S].$ Then Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that $\min\limits_{k=N_{S}+1,\dots,2N_{s}}\|\nu_{S}^{k}\|\leq\varepsilon,$ where $\nu_{S}^{k}\in\partial f({x}_{S}^{k})$ after at most
(26) $\tfrac{16\left[2(p+\nu){C}_{\mathcal{A}}L_{p,\nu}(g)D^{p+\nu-1}\right]^{\frac{1}{p+\nu}}}{\left(1-2^{-1+\frac{1}{p+\nu}}\right)\varepsilon^{\frac{1}{p+\nu}}}=\Theta\left(\tfrac{L_{p,\nu}(g)^{\frac{1}{p+\nu}}D^{\frac{p+\nu-1}{p+\nu}}}{\varepsilon^{\frac{1}{p+\nu}}}\right)$
calls to the $p$-th order oracle.
Proof 3.7.
Recall that the subgradient norm $\|\nu_{S}^{k}\|$ can be decomposed as follows.
$\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|$ $\leq\underbrace{\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|v_{S}^{k}\|}_{\texttt{Term I}}+\underbrace{\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}^{k}-{x}_{i-1}\|^{p+\nu-1}}_{\texttt{Term II}},$
Substituting the convergence rate of the subgradient Term I from 1 and utilizing (17), we have
[…]
$\,\,\,\quad\overset{\text{(i)}}{\leq}{C}_{\mathcal{A}}L_{p,\nu}(g)\left(\tfrac{{C}_{\mathcal{A}}\sigma_{S}D^{p+\nu}}{{{C}_{\mathcal{A}}2^{{p+\nu-2}}(p+\nu)}{4^{(p+\nu)S}}}\right)^{p+\nu-1},$
where in (i), we substitute the linear convergence of $\|x_{s}-x_{s}^{*}\|$ from Lemma 3.4. Furthermore, by the choice of $S$ from (25), we have $\sigma_{S}\geq L_{p,\nu}(g),$ therefore,
$(\texttt{Term I})^{p+\nu}$ $\leq{C}_{\mathcal{A}}\sigma_{S}\left(\tfrac{\sigma_{S}D^{p+\nu}}{{2^{{p+\nu-2}}(p+\nu)}{4^{(p+\nu)S}}}\right)^{p+\nu-1}\leq{C}_{\mathcal{A}}\sigma_{S}\left(\tfrac{\sigma_{S}D^{p+\nu}}{{2^{{p+\nu-1}}}{4^{(p+\nu)S}}}\right)^{p+\nu-1}.$
Substituting the choice for $\sigma_{S}$ (25), we have $\texttt{Term I}\leq\varepsilon/2.$ It remains to bound Term II. By Lemma 3.4, the conditions of Lemma 3.2 holds, thus, we have
Term II $\,\,\,\,\leq\tfrac{2\sigma_{S}D^{p+\nu-1}}{4^{(p+\nu-1)(S-2)}}+\tfrac{(18D)^{p+\nu-1}}{2}\textstyle\sum_{i=1}^{S}\tfrac{\sigma_{i}-\sigma_{i-1}}{4^{i(p+\nu-1)}}$
[…]
Combining the bounds for Term I, II, we have $\|\nu_{s}\|\leq\varepsilon.$ The total calls to the iterations of the subroutines is bounded by
[…]
Similar to Proposition 2.2, the regularization parameter $\sigma_{s}$ starts at $\mathcal{O}(\varepsilon/D^{p+\nu-1})$ and is exponentially increasing, while the epoch length is exponentially decreasing, scaling inversely with respect to the regularization parameter. Algorithm 3 terminates when $\sigma_{S}$ has reached $L_{p,\nu}(g).$ Other choices of the regularization can also be used, with the number of epochs changing accordingly.
In view of Theorem 3.6, we can accelerate the convergence rate of the gradient norm by employing various tensor-based methods as subroutines. Examples include the accelerated cubic Newton method [23], accelerated tensor methods [25], the bi-level unconstrained minimization approach [28], and accelerated proximal-point methods with lower-order solvers [27, 28]. Notably, in [27, 28], they show that second-order methods can achieve the same convergence rate as third-order methods in terms of the function value residual. As a result, using such second-order methods as subroutines in the high-order AR framework Algorithm 3 allows the convergence rate in terms of the gradient norm to match that of the function value residual in [27, 28]. In what follows, we present the Accelerated Regularized Tensor Method (ARTM) [15, 14] as an inner subroutine for solving (3) with $\nu\in[0\,1]$, which improves the gradient complexity bound of ARTM in [15, Theorem 5.7] by removing a logarithmic factor. The gradient complexities for the other methods follow by the same reasoning.
Corollary 3.8.
Suppose the subroutine $\mathcal{A}$ is ARTM. Then, Algorithm 3 can compute an approximate solution $x_{S}^{k}$ s.t., $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within $\mathcal{O}(({D^{{p+\nu-1}{}}}/{\varepsilon})^{\frac{1}{p+\nu}})$ calls to the $p$-th order oracle, where $D$ is defined in (16), and $p\geq 2,$ $\nu\in[0\,\,1].$
Proof 3.9.
[15, Theorem 5.7] shows that 1 holds for ARTM, thus Theorem 3.6 applies.
This result is the tightest known for convex and $p$-times differentiable fucntion with $\nu$-Hölder continuous $p$-th derivative. Although it does not match the lower complexity bounds for such function class  [15, Theorem 6.6], given by $\Omega\left(\left({D^{{p+\nu-1}}}/{\varepsilon}\right)^{\frac{2}{3(p+\nu)-2}}\right).$ The gap arises because no existing methods have yet achieved the optimal rate in terms of the function value residual for such function class. If such methods are developed, the high-order AR framework can directly incorporate them as subroutines $\mathcal{A}$ to accelerate the gradient norm convergence and achieve a matching rate.
3.2 Fast gradient minimization for convex smooth function
In this subsection, we study AR with optimal $p$-th order methods as subroutines to handle composite convex functions with Lipschitz continuous $p$-th derivatives. Specifically, there exists $L_{p+1}(g)>0$ s.t.,
(27) $\|D^{p}g(x)-D^{p}g(y)\|\leq L_{p+1}(g)\|x-y\|,\quad\forall\,\,x,y\in\mathbb{R}^{n}.$
We make the following assumption regarding the subroutine $\mathcal{A}_{s}.$
Assumption 2.
The approximate solution $x_{s}$ exhibits the following performance guarantees: after $N_{s}$ iterations of the subroutine $\mathcal{A}_{s},$ there holds
(28) $f_{s}({x_{s}})-f_{s}(x_{s}^{*})$ $\leq\tfrac{c_{\mathcal{A}}L_{p+1}(g)\|x_{s}^{*}-{x_{s-1}}\|^{p+1}}{N_{s}^{\frac{3p+1}{2}}},\quad\forall\,\,1\leq s\leq S.$
For the $S$-th epoch, there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),N_{S}\leq k\leq 2N_{S},$ such that
(29) $\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|v_{S}^{k}\|\leq\tfrac{{c}_{\mathcal{A}}L_{p+1}(g)\|x_{S}^{*}-x_{S-1}\|^{p}}{{N_{S}^{\frac{3p}{2}}}},$
where $c_{\mathcal{A}}$ is universal constants that depend on $\mathcal{A}_{s}.$
Compared to 1, 2 imposes a faster convergence rate in terms of both the function gap and the gradient norm. Observe that the gradient norm convergence rate still is slower than the function value residual. The above Assumption holds for a wide range of optimal algorithms with different $p.$ The subroutines examples include first order algorithm: the accelerated gradient descent method (AGD) [29], second order algorithm: large stepsize accelerated hybrid proximal extragradient [22], and $p$-th order algorithm: optimal tensor method (OTM) [6, 19, 12] or inexact $p$-th order proximal point method with tensor step [26, 28].
We next show that with properly chosen epoch lengths $\{N_{s}\}_{s\in[S]}$ and 2, high-order AR Algorithm 3 can achieve the linear convergence in terms of distance to the optimal solution across epochs. The proof is similar to Lemma 3.4, thus omitted for simplicity.
Lemma 3.10.
Suppose $p\geq 1,$ and for all $s\in[S],$ the epoch length satisfies
(30) $N_{s}$ $\geq\left\lceil 4\left(\tfrac{(p+1)c_{\mathcal{A}}L_{p+1}(g)}{{\sigma_{s}}}\right)^{\frac{2}{3p+1}}\right\rceil,$
where $D$ is defined in (16). If the subroutine for solving subproblems satisfies 2. Then, for all $s\in[S],$ Algorithm 3 can compute $x_{s}$ such that $\|x_{s}-x_{s}^{*}\|\leq{\|x_{s-1}-x_{s-1}^{*}\|}/{4}.$ Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
With Lemma 3.10 in place, we have the following convergence results of the Algorithm 3, the proof is similar to Theorem 3.6, and thus we omit it for simplicity.
Theorem 3.11.
Suppose the Assumptions in Lemma 3.10, and
(31) $S$ $\coloneqq\left\lceil\tfrac{2}{3p+1}\log_{2}\tfrac{{{C}_{\mathcal{A}}D^{p}L_{p+1}(g)}}{\varepsilon}\right\rceil+1,\quad\sigma_{s}\coloneqq\tfrac{2^{\frac{(s-1)(3p+1)}{2}}\varepsilon}{{C}_{\mathcal{A}}D^{p}},$
for all $s\in[S].$ Then Algorithm 3 can compute an approximate solution $x_{S}^{k}$ together with $\nu_{S}^{k},$ such that $\nu_{S}^{k}\in\partial f(x_{S}^{k}),$ and $\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$ after at most
(32) $\tfrac{8\left[(p+1)L_{p+1}(g)D^{p}\right]^{\frac{2}{3p+1}}}{\varepsilon^{\frac{2}{3p+1}}}=\Theta\left(\left[\tfrac{L_{p+1}(g)D^{p}}{\varepsilon}\right]^{\frac{2}{3p+1}}\right)$
calls to the $p$-th order oracle.
We can utilize Theorem 3.11 to derive optimal algorithms in terms of (sub)gradient norm minimization for convex composite problem (3) by employing various tensor-based methods as subroutines. We illustrate with the following examples.
Corollary 3.12.
Suppose the subroutine $\mathcal{A}$ is A-NPE [22]. Then, a third-order AR framework Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within
$\mathcal{O}([{L_{3}(g)D^{p}}/{\varepsilon}]^{\frac{2}{7}}\log\tfrac{1}{\varepsilon})$
evaluations of the gradient and Hessian, where $D$ is defined in (16).
Proof 3.13.
[22, Theorem 4.1] shows that 2 holds for A-NPE, thus Theorem 3.11 applies.
We next show a gradient norm minimization using a second-order method for a third-order smooth problem, and show its convergence rate of the gradient norm matches that of the function value residual in [26]. For simplicity, we consider the unconstrained problem (1) in this example.
Corollary 3.14.
Suppose the subroutine $\mathcal{A}$ is the second order implementation of the inexact third-order proximal point method [26]. Then, a fourth-order AR Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that
$\min_{k=N_{S}+1,\dots,2N_{S}}\|\nabla f(x_{S}^{k})\|\leq\varepsilon$
within $\mathcal{O}([{L_{4}(f)D^{3}}/{\varepsilon}]^{\frac{1}{5}})$ iterations, where $D$ is defined in (16), at each iteration, it calls the second-order and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
Proof 3.15.
[26, Theorem 3.8, 5.5] show that 2 holds with $p=3$, thus Theorem 3.11 applies.
4 Parameter-free algorithm for convex problems
For all convergence results in the Section 2 and Section 3 are based on the exact solutions of the auxiliary optimization problem at each iteration of the high-order method, and assumed to have access to certain key problem parameters, including the Lipschitz constant of the $\nu$-Hölder continuous $p$-th derivative $L_{p,\nu}(g)$, and an upper bound $D$ on the initial distance to the solution set, such that $\text{dist}(x_{0},X^{*})\leq D$. However, in generally, auxiliary optimization problem at each iteration can not be solved in general, and those problem dependent parameters are not known before hand. In this section, we investigate inexact and parameter-free implementations of high-order methods for gradient minimization that do not rely on such information. For clarity of exposition, we first focus on the unconstrained setting; analogous parameter-free strategies can be developed for problems with simple constraints composite structures.
Algorithm 4 High-order AR without the knowledge of $L_{p,\nu}$
1: Initial regularization $\sigma_{0}=0,\sigma_{1}>0,$ and initial point $x_{0}\in\mathbb{R}^{n},$ and $L_{0}.$
2: for $s=1,2\dots$ do
3:   If $s>1,$ set $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1}.$
4:   Compute an approximate solution $x_{s}$ of the proximal subproblem (33) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$
5:   where $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2$ by running $\mathcal{A}_{s}$ with the initialization $x_{s-1}$ and initial line search value $L_{s-1}\coloneqq L_{{s-1},{N_{s-1}}}.$ $\mathcal{A}_{s}$ will choose $N_{s}$ as $k,$ and output the $k$-th iterate as $x_{s},$ its line search value $L_{s,k}$ when $k\geq 8\left[\tfrac{L_{{s},{k}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1.$ Continue the same number of surplus iterations to record $L_{s,{2N_{s}}}.$
6:   If $\sigma_{s}\geq\tfrac{(L_{s,{2N_{s}}})^{(p+\nu)^{2}}}{(L_{s,{N_{s}}})^{(p+\nu-1)(p+\nu+1)}},$ then terminate with $\widehat{x}=\operatorname*{arg\,min\,}\limits_{N_{s}<k\leq 2N_{s}}\{\|\nabla f(z_{s}^{k})\|\}.$
7: output $\widehat{x}$
Denote $(s,k)$ as the pair referring to the $k$-th iteration of epoch $s,$ and $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$ We make the following Assumption on the subroutine $\mathcal{A}_{s}$ used to solve (33).
Assumption 3.
The subroutine $\mathcal{A}_{s}$ has the following performance guarantee: for any $k>1,$ the following bound holds
(34) $f_{s}({x_{s}^{k}})-f_{s}(x_{s}^{*})$ $\leq\tfrac{L_{s,k}\|{x_{s-1}}-x_{s}^{*}\|^{p+\nu}}{(k-1)^{p+\nu}},\quad\forall\,\,1\leq s\leq S,$
where $x_{s}^{k}$ is the computed approximate solution and $L_{s,k}$ is an local estimate of Hölder constant of $f$ at $(s,k)$ such that
(35) $L_{s,k}\leq c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\},$
where $c_{\mathcal{A}}$ is a universal constant depending on the subroutine $\mathcal{A}_{s},$ and $\theta$ is a user-defined subproblem inexactness parameter used in $\mathcal{A}_{s}.$ For each $s$-th epoch, there exists sequences $\{{z}_{s}^{k}\},$ such that $f({z}_{s}^{k})\leq f(x_{s}^{k})$ and $\nabla f_{s}({z}_{s}^{k}),N_{s}\leq k\leq 2N_{s},$ such that
(36) $\min\limits_{k=N_{s}+1,\dots,2N_{s}}\|\nabla f_{s}({z}_{s}^{k})\|\leq\tfrac{L_{s,2N_{s}}\|x_{s-1}-x_{s}^{*}\|^{p+\nu-1}}{(2N_{s}-1)^{\frac{(p+\nu-1)(p+\nu+1)}{p+\nu}}},$
where $L_{s,2N_{s}}$ is the local estimate of Hölder constant at $(s,2N_{s})$ such that
(37) $L_{s,{2N_{s}}}\leq c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}.$
Notice that 3 relaxes 1 in terms of two aspects. First, it does not require the knowlwdge of $L_{p,\nu}(f),$ the $\nu$-Hölder continuity constant of the $p$-th order derivatives. Instead, it only requires a local estimate $L_{s,k}.$ Second, it allows each iteration of the subroutine $\mathcal{A}_{s}$ to be solved inexactly; this inexactness can be characterized by a parameter $\theta,$ where smaller $\theta$ means a more accurate solution for each iteration. This assumption can be easily satisfied by many algorithms, for example, line search based accelerated tensor methods [14, 15, 16].
Theorem 4.1.
Suppose the subroutine satisfies 3. Given $x_{0},L_{0}$ and $\sigma_{1},$ and suppose $\sigma_{1}\leq c_{\mathcal{A}}.$ Furthermore, choose the regularization parameter $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1},$ for all $s\geq 2,$ and the epoch length $N_{s}$ as
(38) $N_{s}\geq 8\left[\tfrac{L_{s,N_{s}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1,\forall s\geq 1.$
Then Algorithm 4 can compute an approximate solution $\widehat{x}$ such that
(39) $\|\nabla f(\widehat{x})\|\leq 3\sigma_{1}[9\cdot\textnormal{dist}(x_{0},X^{*})]^{p+\nu-1}$
after $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}}{\sigma_{1}}\right]^{\frac{1}{p+\nu}}\right)$ calls to the $p$-th order oracle.
Proof 4.2.
Observe that the gradient norm $\|\nabla f(z_{S}^{k})\|$ can be decomposed as follows.
$\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|\nabla f(z_{S}^{k})\|$
$\leq\underbrace{\min\limits_{k=N_{S}+1,\dots,2N_{S}}\|\nabla f_{S}(z_{S}^{k})\|}_{\texttt{Term I}}+\underbrace{\textstyle\sum_{i=1}^{S}(\sigma_{i}-\sigma_{i-1})\|x_{S}^{k}-{x}_{i-1}\|^{p+\nu-1}}_{\texttt{Term II}},$
We start with bounding Term II. By the uniform convexity of $f_{s}$ and the optimality condition $\nabla f_{s}(x_{s}^{*})=0,$ we have
(40) $f_{s}(x_{s})-f_{s}(x_{s}^{*})$ $\geq\tfrac{\sigma_{s}}{p+\nu}\left(\tfrac{1}{2}\right)^{p+\nu-2}\|x_{s}-x_{s}^{*}\|^{p+\nu},\quad\forall s\in[S].$
Similar to the proof of Proposition 2.2, by using the optimality condition of (33), for all $s\geq 1,$ $s\in[S],$ there holds
(41) $\|x_{s-1}-x_{s}^{*}\|$ $\leq\|x_{s-1}-x_{s-1}^{*}\|.$
Combining it with (34) in 3, we have
[…]
Combining it with the subroutine sample size condition (38), we have
(43) $\|{x_{s}}-x_{s}^{*}\|\leq\tfrac{1}{4}\|{x_{s-1}}-x_{s-1}^{*}\|.$
Similarly, notice that in each epoch $s$, there holds
$\tfrac{\sigma_{s}}{p+\nu}\left(\tfrac{1}{2}\right)^{p+\nu-2}\|z_{s}^{k}-x_{s}^{*}\|^{p+\nu}$ $\,\,\leq f_{s}(z_{s}^{k})-f_{s}(x_{s}^{*})$
[…]
Combining it with the subroutine sample size condition (38), we have $\|{z_{s}^{k}}-x_{s}^{*}\|\leq\tfrac{1}{4}\|{x_{s-1}}-x_{s-1}^{*}\|.$ Therefore, the conditions of Lemma 3.2 holds, thus, we have
Term II $\,\,\,\,\leq\tfrac{2\sigma_{S}\|x_{0}-x^{*}\|^{p+\nu-1}}{4^{(p+\nu-1)(S-2)}}+\tfrac{(18\|x_{0}-x^{*}\|)^{p+\nu-1}}{2}\textstyle\sum_{i=1}^{S}\tfrac{\sigma_{i}-\sigma_{i-1}}{4^{i(p+\nu-1)}}$
$\,\,\overset{\text{(i)}}{\leq}\tfrac{2\cdot 2^{(p+\nu-1)(S-1)}\sigma_{1}\|x_{0}-x^{*}\|^{p+\nu-1}}{4^{(p+\nu-1)(S-2)}}+\tfrac{(18\|x_{0}-x^{*}\|)^{p+\nu-1}}{2}\textstyle\sum_{i=1}^{S}\tfrac{2^{(p+\nu-1)(i-1)}\sigma_{1}}{4^{i(p+\nu-1)}}$
$\,\,\overset{\text{(ii)}}{\leq}2\sigma_{1}(9\|x_{0}-x^{*}\|)^{p+\nu-1},$
where $x^{*}$ is any solution in the optimal solution set $X^{*}.$ In (i), we substitute the choice for $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1}.$ In (ii), we used $p+\nu\geq 2,$ and $S\geq 1.$ It remains to bound Term I. Substituting the convergence rate of the slow gradient norm from 3 and (41), we have
[…]
$\,\,\,\,\overset{\text{(iii)}}{\leq}\sigma_{S}\left(\tfrac{{16\sigma_{S}\|x_{S-1}^{*}-x_{S-1}\|^{p+\nu}}{}}{16^{p+\nu}(p+\nu)}\right)^{p+\nu-1}\overset{\text{(iv)}}{\leq}\sigma_{S}\left(\tfrac{16\sigma_{S}\|x_{0}-x^{*}\|^{p+\nu}}{{(p+\nu)}{4^{(p+\nu)(S+1)}}}\right)^{p+\nu-1},$
where in (iii), we used the termination criterion for (4). In (iv), we substitute the linear convergence of $\|x_{s}-x_{s}^{*}\|$ from (43). Substituting the choice of $\sigma_{S}=2^{(p+\nu-1)(S-1)}\sigma_{1},$ we have
(44) Term I $\leq 2^{(p+\nu-1)(S-1)}\sigma_{1}\left(\tfrac{16\|x_{0}-x^{*}\|^{p+\nu}}{{(p+\nu)}{4^{(p+\nu)(S+1)}}}\right)^{\frac{p+\nu-1}{p+\nu}}\leq\sigma_{1}\|x_{0}-x^{*}\|^{p+\nu-1}.$
Combining the errors from Term I, II, we have
[…]
$\leq 2\sigma_{1}(9\|x_{0}-x^{*}\|)^{p+\nu-1}+\sigma_{1}\|x_{0}-x^{*}\|^{p+\nu-1}\leq 3\sigma_{1}(9\|x_{0}-x^{*}\|)^{p+\nu-1}.$
Furthermore, by the termination criterion (4), we have
[…]
Therefore, we have
$S$ $\leq 2+\tfrac{(p+\nu)^{2}}{{p+\nu-1}}\log_{2}\left(\tfrac{c_{\mathcal{A}}^{\frac{1}{(p+\nu)^{2}}}\max\{pL_{p,\nu}(f),L_{0},\theta\}}{\sigma_{1}^{\frac{1}{(p+\nu)^{2}}}}\right)-(p+\nu+1)\log_{2}L_{0}.$
Observe that the RHS is positive due to the condition $\sigma_{1}\leq c_{\mathcal{A}}.$ Note that such a condition is easy to satisfy as one is interested in smaller precision $\sigma_{1},$ and $c_{\mathcal{A}}$ is a constant related to the subroutine $\mathcal{A}.$ The total calls to oracle reads
[…]
$\,\,\leq 32\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)}{4\sigma_{1}}\right]^{\frac{1}{p+\nu}}+S.$
This concludes the proof.
The expression (39) sheds light on the impact of the initial regularization parameter $\sigma_{1}$ on the convergence. Specifically, the following comments are in order.
  • 1)
    Matching the convergence rate of the function gap residual: If we choose $\sigma_{1}={\varepsilon}/{\{3[9\cdot\text{dist}(x_{0},X^{*})]^{p+\nu-1}\}},$ then Theorem 4.1 shows that Algorithm 4 can compute an approximate solution $\widehat{x}$ s.t. $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    (46) $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}\text{dist}(x_{0},X^{*})^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th order oracle. This convergence result for the inexact and adaptive gradient minimization is new, and it matched its faster convergence rate for function value residual used in the subroutine $\mathcal{A}$ (cf. 3), see for example, [14, 15, 18, 16]. It improves over the current convergence result for adaptive and inexact gradient minimization as (36) shown in [16].
  • 2)
    Unknown initial optimality gap $D$: However, in general $\text{dist}(x_{0},X^{*})$ is unknown, and we may use $D$ to estimate it. If $D>\text{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ we can compute an approximate solution $\widehat{x}$ using Algorithm 4 such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}D^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th order oracle, which is worse than (46). Next we show that we can avoid such case by choosing $D$ properly. Specifically, define $D^{p+\nu-1}\coloneqq\|\nabla f_{1}({x}_{1}^{2})\|/L_{1,2},$ by (36) with $s=1,N_{1}=1,$ we have
[…]
    where $x^{*}$ is any optimal solution in $X^{*}.$ Therefore, we have $D\leq\text{dist}(x_{0},X^{*}).$ If $D<\text{dist}(x_{0},X^{*}),$ by choosing $\sigma_{1}={\varepsilon}/[3(9D)^{p+\nu-1}],$ the computed approximate solution $\widehat{x}$ only satisfies $\|\nabla f(\widehat{x})\|\leq\varepsilon[\text{dist}(x_{0},X^{*})/{D}]^{p+\nu-1}.$ Therefore, it does not attain the desired accuracy $\varepsilon.$ In this case, we simply discard all previous results and restart the computation with a larger guess $D,$ such a procedure that gradually increases $D$ is called a “guess-and-check” procedure. It is first used parameter-free optimal gradient minimization using first order method [21]. We will show below that such a strategy maintains the same oracle complexity as if $\operatorname{dist}(x_{0},X^{*})$ were known, while still guaranteeing $\|\nabla f(\widehat{x})\|\leq\varepsilon$.
Algorithm 5 A guess-and-check implementation of Algorithm 4
1: Initial $\sigma_{0}=0,$ $x_{0}\in\mathbb{R}^{n},$ and $L_{0}.$ Target accuracy $\varepsilon.$
2: Run the subroutine $\mathcal{A}_{s}$ with $s=1,N_{1}=1,$ define $D_{0}^{p+\nu-1}\coloneqq\|\nabla f_{1}({x}_{1}^{2})\|/L_{1,2}.$
3: for $t=1,2\dots$ do
4:   Set $D_{t}=4D_{t-1}.$
5:   Compute $\widehat{x}=\text{AR}(f,x_{0},{\varepsilon}/{[3(9D_{t})^{p+\nu-1}]}).$
6:   If $\|\nabla f(\widehat{x})\|\leq\varepsilon,$ then terminate with $\widehat{x}$.
We have the following convergence guarantee for Algorithm 5.
Theorem 4.3.
Suppose the Assumptions in Theorem 4.1. Then Algorithm 5 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most (46) calls to the $p$-th order oracle.
Proof 4.4.
Observe that after at most $T$ calls to Algorithm 4, Algorithm 5 will terminate, where $T$ satisfies $D_{T}=4^{T}D_{0}\geq\text{dist}(x_{0},X^{*}),$ and $D_{T-1}=4^{T-1}D_{0}\leq\text{dist}(x_{0},X^{*}),$ therefore, we have $T-1\leq\lceil\log_{4}\tfrac{\text{dist}(x_{0},X^{*})}{D_{0}}\rceil.$ By Theorem 4.1, in total, Algorithm 5 requires
$c\textstyle\sum_{t=1}^{T-1}\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0}\cdot 4^{t})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9D_{0})^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{4^{\frac{(T-1)(p+\nu-1)}{p+\nu}}-1}{4^{\frac{p+\nu-1}{p+\nu}}-1}$
$\leq c\left[\tfrac{c_{\mathcal{A}}\max\{pL_{p,\nu}(f),L_{0},\theta\}(p+\nu)[3(9\text{dist}(x_{0},X^{*}))^{p+\nu-1}]}{4{\varepsilon}}\right]^{\frac{1}{p+\nu}}\tfrac{1}{4^{\frac{p+\nu-1}{p+\nu}}-1},$
calls to the $p$-th order oracle, where $c$ is a universal constant. This concludes the proof.
5 Gradient Minimization Framework for Uniformly Convex Problems
In this section, we show that by using optimal (sub)gradient norm convergence results of the high-order AR framework for convex composite problems, we can obtain new fast (sub)gradient norm convergence rates for composite problems satisfying a certain regularity condition.
5.1 Gradient minimization for convex composite problems
Consider a class of composite problems (3) and in addition, $f$ is uniformly convex on $\mathbb{R}^{n}$ of degree $q$ with parameter $\sigma_{q}(f)>0$ [31, 4, 33, 35], i.e.,
(47) $f(x)\geq f(y)+\langle g_{y},x-y\rangle+\tfrac{\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,x,y\in\mathbb{R}^{n},g_{y}\in\partial f(y).$
Consider restarting the high-order AR framework Algorithm 3 whenever the (sub)gradient norm is reduced by half. Specifically, in each epoch $k,$ we run $m_{k}$ iterations of high-order AR framework Algorithm 3 with the initialization $x_{k-1},$ i.e., the output of the previous epoch k1.k-1. By a careful choice of the epoch length $m_{k}$, which depends on the relation between the upper smoothness level $p$ and the lower uniform convexity level $q$, we can ensure the existence of a subgradient sequence with exponentially decaying norms. We present below the high-order AR framework for uniformly convex composite problems together with its convergence guarantee.
Algorithm 6 Uniformly convex accumulative regularization
1: Initial point $x_{0}\in\mathbb{R}^{n},$ $L_{p+1}(g),\sigma_{q}(f).$
2: for $k=1,2\dots,$ do
3:   Compute an approximate solution $(x_{k},\nu_{k})$ of $f$ with initialization $(x_{k-1},\nu_{k-1})$ by running high-order AR framework in Algorithm 3 for $m_{k}$ iterations, where $m_{k}=\max\left\{\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left({\tfrac{q\|\nu_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{2(p-q+1)}{(3p+1)(q-1)}},1\right\}.$
4:   If $\|\nu_{k}\|\leq\varepsilon,$ then terminate with $\widehat{x}=x_{k}.$
5: output $\widehat{x}$
Proposition 5.1.
Suppose that $\|\nu_{0}\|\geq\varepsilon.$ Then, Algorithm 6 can compute a point $\widehat{x}$ such that $\|\widehat{\nu}\|\leq\varepsilon$ within at most
$\mathcal{O}\left(\left[\tfrac{(p+1)L_{p+1}(g)}{{\sigma_{q}(f)}}\right]^{\frac{2}{3p+1}}\log\tfrac{\|\nu_{0}\|}{\varepsilon}\right)$
calls to the $p$-th order oracle if $q=p+1;$ and at most
$\mathcal{O}\left(\left[\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right]^{\frac{2}{3p+1}}\left(\tfrac{q}{\sigma_{q}(f)}\right)^{{\frac{2(q-1-p)}{(3p+1)(q-1)}}}\left(\tfrac{1}{\varepsilon}\right)^{{\frac{2(q-1-p)}{(3p+1)(q-1)}}}\right)$
calls to the $p$-th order oracle if $q>p+1;$ and at most
$\mathcal{O}\left(\tfrac{1}{\log{\left(\frac{p}{q-1}\right)}}{\log\log\left[\tfrac{\sigma_{q}(f)}{{q\varepsilon}}{\left(\tfrac{\sigma_{q}(f)}{qL_{p+1}(g)}\right)^{\frac{q-1}{p+1-q}}}\right]}+\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}\right)$
calls to the $p$-th order oracle if $q<p+1.$
Proof 5.2.
By adding two copies of inequality (47), we have
(48) $\langle x-y,g_{x}-g_{y}\rangle\geq\tfrac{2\sigma_{q}(f)}{q}\|x-y\|^{q},\quad\forall\,\,x,y\in\mathbb{R}^{n},$
and for all $g_{x}\in\partial f(x),g_{y}\in\partial f(y).$ Substituting $x=x_{k-1}$ and $y=x^{*}$ into (48), and using the Cauchy-Swartcz inequality and optimality condition $0\in\partial f(x^{*}),$ we can choose $g_{x^{*}}=0$ and thus, we have
(49) $\|x_{k-1}-x^{*}\|\leq\left({\tfrac{q\|g_{k-1}\|}{2\sigma_{q}(f)}}\right)^{\frac{1}{q-1}},$
where $g_{k-1}$ is an arbitrary sub-gradient of $f$ at point $x_{k-1}.$ Let us prove by induction that there exists a sequence $\{\nu_{k}\}_{k\geq 1}$ such that $\nu_{k}\in\partial f(x_{k})$ and $\|\nu_{k}\|\leq\|\nu_{0}\|/2^{k}.$ Suppose it is true for k1.k-1. At epoch $k,$ we restart the high-order AR framework Algorithm 3 with the initial point $x_{k-1}.$ By Theorem 3.11, Algorithm 3 can compute an approximate solution $x_{k}$ together with $\nu_{k}$ such that
[…]
Given that $g_{k-1}$ can be any sub-gradient of $f$ at point $x_{k-1},$ then we choose $g_{k-1}=\nu_{k-1}.$ By the choice of $m_{k}$ in Algorithm 6, we have $\|\nu_{k}\|\leq{{\|\nu_{k-1}\|}/{2}}\leq\|\nu_{0}\|/2^{k}.$ By simply restarting the high-order AR framework whenever the norm of the approximate solution $\nu_{k}$ is reduced by half, we have the following convergence guarantees.
Case I: $q=p+1.$ In this case, $m_{k}\equiv m_{0}=\left[{qL_{p+1}(g)}/{{\sigma_{p+1}(f)}}\right]^{\frac{2}{3p+1}},$ thus the convergence rate is linear. We can obtain an approximate solution $\widehat{x}$ such that $\|\hat{\nu}\|\leq\varepsilon$ within at most $\mathcal{O}(\log{\|\nu_{0}\|}/{\varepsilon})$ epochs.
Case II: $q>p+1.$ In this case, $m_{k}$ increases exponentially, to reach a point ${x}_{K}$ such that $\|\nu_{K}\|\leq\varepsilon,$ we need
$\textstyle\sum_{k=1}^{K}m_{k}$ $=\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{\sigma_{q}(f)}{q\|\nu_{0}\|}\right)^{\frac{2(q-p-1)}{(3p+1)(q-1)}}\textstyle\sum_{k=1}^{K}2^{\frac{2(q-p-1)k}{(3p+1)(q-1)}}$
$\overset{\text{(a)}}{\leq}\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{2\sigma_{q}(f)}{q}\right)^{{\frac{2(q-p-1)}{(3p+1)(q-1)}}}\tfrac{1}{2^{\frac{2(q-p-1)}{(3p+1)(q-1)}}-1}\left(\tfrac{1}{\varepsilon}\right)^{{\frac{2(q-p-1)}{(3p+1)(q-1)}}}$
calls to the $p$-th order oracle, where in (a), we used $\|\nu_{K}\|\leq\|\nu_{0}\|/2^{K}\leq\varepsilon.$
Case III: $q<p+1.$ In this case, $m_{k}$ decays exponentially until it reaches $1,$ then, the algorithm converges superlinearly. Thus, the algorithm has two phases. Let $k=K_{1}$ denote the final iteration count of Phase I, where $m_{K_{1}}=1$ and $m_{K_{1}-1}\geq 1,$ we have
(50) $\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{K_{1}-1}\|}{2\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}\leq 1,\quad\text{and}\quad m_{K_{1}}=1.$
At epoch $K_{1},$ by Theorem 3.11, we have
$\|\nu_{K_{1}}\|$ $\,\,\,\,\leq\tfrac{L_{p+1}(g)\|x_{K_{1}-1}-x^{*}\|^{p}}{m_{K_{1}}^{\frac{3p+1}{2}}}={L_{p+1}(g)\|x_{K_{1}-1}-x^{*}\|^{p}}$
[…]
Define $H=\left(\tfrac{1}{L_{p+1}(g)}\right)^{\frac{q-1}{p+1-q}}\left(\tfrac{2\sigma_{q}(f)}{q}\right)^{\frac{p}{p+1-q}},$ then we have $\tfrac{\|\nu_{K_{1}}\|}{H}\leq\tfrac{1}{2}.$ The total iterations of Phase I reads
$\textstyle\sum_{k=1}^{K_{1}}m_{k}$ $\leq\textstyle\sum_{k=1}^{K_{1}}\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{2^{k}\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}$
$\leq c\left(\tfrac{qL_{p+1}(g)}{{\sigma_{q}(f)}}\right)^{\frac{2}{3p+1}}\left(\tfrac{q\|\nu_{0}\|}{\sigma_{q}(f)}\right)^{{\frac{2(p-q+1)}{(3p+1)(q-1)}}}.$
Then, Algorithm 6 enters into Phase II: superlinear convergence phase. For all $k\geq K_{1},$ $m_{k}=1,$ by Theorem 3.11, $\|\nu_{k}\|\leq L_{p+1}(g)\left({{q\|\nu_{k-1}\|}/{2\sigma_{q}(f)}}\right)^{\frac{p}{q-1}},$ thus ${\|\nu_{k}\|}/{H}\leq\left({\|\nu_{k-1}\|}/{H}\right)^{\frac{p}{q-1}}.$ Therefore, for all $k\geq K_{1},$ we have $\|\nu_{k}\|\leq H\left(\tfrac{1}{2}\right)^{(\frac{p}{q-1})^{k-K_{1}}}.$ For any $\varepsilon>0,$ to reach a point $\widehat{x}$ to satisfy $\|\widehat{\nu}\|\leq\varepsilon,$ we need $k\geq\tfrac{\log\log\tfrac{H}{\varepsilon}-\log\log 2}{\log{\left(\frac{p}{q-1}\right)}}+K_{1}$ iterations.
Depending on the relation between $p+1$ and $q,$ Algorithm 6 can exhibit linear or superlinear convergence rate. Such complexities in terms of the gradient norm have not been derived before.
5.2 parameter-free gradient minimization for uniformly convex function
In this section, we develop parameter-free and inexact method for uniformly convex function. In particular, we do not assume the knowledge of Lipschitz continuous parameter and uniform convex parameter. We will utilize the parameter-free gradient minimization methods developed from Section 4, and use a guess and search procedure to estimate the uniform convex parameter.
For simplicity, we will focus on the unconstrained problem (1) where $f$ is $p$-times differentiable with a Lipschitz continuous $p$-th derivative (2), and furthermore $f$ is uniformly convex with order $q=p+1.$ We have the following parameter-free and inexact algorithm for uniformly convex function.
Algorithm 7 A parameter-free uniformly convex accumulative regularization
1: Initial point $f,\varepsilon,x_{0},\sigma_{p+1,0},L_{0}.$
2: for $t=1,2\dots,$ do
3:   Set $(x_{t},L_{t})=\text{AR}(f,x_{t-1},\tfrac{\sigma_{p+1,t-1}}{3(p+1)9^{p}},L_{t-1}).$
4:   If $\|\nabla f(x_{t})\|>\tfrac{\|\nabla f(x_{t-1})\|}{2},$ then $\sigma_{p+1,t}=\tfrac{\sigma_{p+1,t-1}}{4}.$
5:   If $\|\nabla f(x_{t})\|\leq\varepsilon,$ then terminate with $\widehat{x}=x_{t}.$
6: output $\widehat{x}$.
Theorem 5.3.
Suppose the Assumptions in Theorem 4.1 and $f$ satisfies the uniform convexity with parameter $\sigma_{p+1}(f)$ of order $p+1.$ Then Algorithm 7 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right)$
calls to the $p$-th order oracle.
Proof 5.4.
i) By Theorem 4.1 with $\nu=1,$ for each epoch $t,$ we have
[…]
after $\mathcal{O}([{\max\{{pL_{p+1}}(f),L_{0},\theta\}}/{\sigma_{p+1,t-1}}]^{\frac{1}{p+1}})$ calls to the $p$-th order oracle.
ii) Algorithm 7 terminates after
$t\geq T=\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil+\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$
calls to Algorithm 4. This is because if we have $\|\nabla f(x_{t})\|>{\|\nabla f(x_{t-1})\|}/{2},$ then by (51), we conclude the current $\sigma_{p+1,t-1}$ overestimate the true uniform convexity parameter $\sigma_{p+1}(f).$ Thus we set $\sigma_{p+1,t}={{\sigma_{p+1,t-1}}}/{{4}}.$ However the estimate $\sigma_{p+1,t}$ will no longer change after at most $T_{1}=\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil$ epochs.
On the other hand, if we have $\|\nabla f(x_{t})\|\leq{\|\nabla f(x_{t-1})\|}/{2},$ then by (51), we conclude the current $\sigma_{p+1,t-1}$ is smaller enough, and thus do not need to decrease. Furthermore, $\|\nabla f(x_{t})\|\leq\varepsilon$ after at most $T_{2}=\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$ epochs.
Combining i), ii), the total calls to the $p$-th order oracle reads
$\textstyle\sum_{t=1}^{T}c_{p}\left[\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1,t-1}}\right]^{\frac{1}{p+1}}=c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}\textstyle\sum_{t=1}^{T}\left(\tfrac{1}{\sigma_{p+1,t-1}}\right)^{\frac{1}{p+1}}$
$\overset{\text{(i)}}{\leq}c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}\textstyle\sum_{t=1}^{T_{1}}\left(\tfrac{4^{t-1}}{\sigma_{p+1,0}}\right)^{\frac{1}{p+1}}+\tfrac{c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}T_{2}}{\sigma_{p+1}(f)^{\frac{1}{p+1}}}$
$\overset{\text{(ii)}}{\leq}c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\tfrac{4^{\frac{1}{p+1}}}{4^{\frac{1}{p+1}}-1}+c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil,$
where $c_{p}=3(p+1)9^{p},$ and in (i), we used the fact $\sigma_{p+1,t}=\sigma_{p+1,t^{\prime}}$ for all $T_{1}\leq t^{\prime}\leq T,$ and $\sigma_{p+1,T_{1}}=\sigma_{p+1,0}/{4^{T_{1}}}\geq\sigma_{p+1}(f);$ and in (ii) we used the definition of $T_{1}$ and $T_{2}.$
Observe the convergence rate of Theorem 5.3 matches the convergence rate when assuming $\sigma_{p+1}(f)$ is known. Similar guess and check procedure for unknown $\sigma_{p+1}(f)$ can easily extended to when $p+1\neq q,$ and thus omitted for simplicity.
6 Concluding Remarks
In this paper, we establish a unified framework of high-order accumulative regularization for gradient norm minimization over convex and uniformly convex functions. By carefully designing regularization strategies and leveraging high-order information, the proposed algorithms bridge the long-standing gap between function-residual and gradient-based complexity guarantees. Our results demonstrate that the high order framework effectively accelerates the existing slow convergence of the gradient norm by exploiting the subroutine’s fast function-residual decrease. Furthermore, we develop several parameter-free variants that achieve the same convergence rates without requiring prior knowledge of problem parameters. Overall, our framework attains best-known gradient norm rates across a broad range of smoothness and curvature regimes, recovering classical first-order results as special cases while extending to previously unexplored high-order settings.
Appendix A Proof of Lemma 2.1
Proof A.1.
We aim to maintain the following relations $\mathcal{R}_{k}^{1},\mathcal{R}_{k}^{2}$ for all $k\geq 1.$ We prove by induction. When $k=1,$ by the definition of $f_{1}(x)$ in Algorithm 1, we have $f_{1}^{*}=f(x_{1})+\sqrt{1/{(L+M)}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}.$ By choosing $A_{1}=1,$ we conclude $\mathcal{R}_{1}^{1}$ holds. Furthermore, we have
$f_{1}(x)$ $\,=f(x_{1})+\tfrac{1}{\sqrt{L+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\overset{\text{(a)}}{\leq}\min\limits_{y\in\mathbb{R}^{n}}\left[f(y)+\tfrac{2L\|y-x_{0}\|^{3}}{6}\right]+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\,\leq f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}},$
where in (a), we used [30, Lemma 4], [23, Lemma 6], and in the last step, we choose $y=x.$ Thus, $\mathcal{R}_{1}^{2}$ holds. Assume that $\mathcal{R}_{k}^{1}$ and $\mathcal{R}_{k}^{2}$ hold for for some $k\geq 1.$ For $k+1,$
$f_{k+1}(x)\,$ $\,=f_{k}(x)+a_{k}\left[f(x_{k+1})+\langle\nabla f(x_{k+1}),x-x_{k+1}\rangle\right]$
$\overset{\mathcal{R}_{k}^{2}}{\leq}A_{k}f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}$
$\quad+a_{k}[f(x_{k+1})+\langle\nabla f(x_{k+1}),x-x_{k+1}\rangle]$
$\overset{\text{(b)}}{\leq}A_{k+1}f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}},$
where in (b), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ Therefore $\mathcal{R}_{k+1}^{2}$ holds. It remains to show $\mathcal{R}_{k+1}^{1}$ holds, which it is a straightforward modification of [23], thus omitted for simplicity. By the definition of $f_{k+1}(x),$ we have
(52) $f_{k+1}^{*}$ $\,\,{=}\min\limits_{x\in\mathbb{R}^{n}}\left\{f_{k}(x)+a_{k}\left[f(x_{k+1})+\langle\nabla f(x_{k+1}),x-x_{k+1}\rangle\right]\right\}.$
We proceed with providing a lower bound on $f_{k}(x).$ By the definition of $f_{k}(x)$ in Algorithm 1, we have $f_{k}(x)=\ell(x)+\textstyle\sum_{i=1}^{k-1}a_{i}\left(f(x_{i+1})+\langle\nabla f(x_{i+1}),x-x_{i+1}\rangle\right)+{C}\|x-x_{0}\|^{3}/6,$ therefore, $f_{k}(x)$ is an uniform strongly convex function. Thus, we have
(53) $f_{k}(x)\overset{\text{(b)}}{\geq}$ $f_{k}^{*}+\tfrac{C}{12}\|x-\nu_{k}\|^{3}\ \overset{\mathcal{R}_{k}^{1}}{\geq}A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}+\tfrac{C\|x-\nu_{k}\|^{3}}{12},$
where in (b), we used the uniform strong convexity of $f_{k}$ $\nu_{k}$ is the solution of $f_{k}(x)$ defined in step 2 of Algorithm 1, and $\nabla f_{k}(\nu_{k})=0.$ Substituting the lower bound for $f_{k}(x)$ into (52), we have
[…]
$\quad\left.+a_{k}\left[f(x_{k+1})+\langle\nabla f(x_{k+1}),x-x_{k+1}\rangle\right]\right\}$
$\overset{\text{(a)}}{\geq}A_{k+1}[f(x_{k+1})+\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle]+\textstyle\sum_{j=1}^{k}{\tfrac{A_{j}\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}$
(54) $\quad+\min\limits_{x\in\mathbb{R}^{n}}\left\{a_{k}\langle\nabla f(x_{k+1}),x-\nu_{k}\rangle+\tfrac{C}{12}\|x-\nu_{k}\|^{3}\right\},$
where in (a), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ By [30, Lemma 4], [23, Lemma 6], if $M\geq 2L,$ there holds
(55) $\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle\geq\sqrt{\tfrac{2}{L+M}}\|\nabla f(x_{k+1})\|^{\frac{3}{2}}.$
Furthermore, given that $a_{k},C,M$ are chosen as in Lemma 2.1, we have
$A_{k}$ $=A_{k-1}+a_{k}=\tfrac{k(k+1)(k+2)}{6},\quad a_{k}^{-\frac{3}{2}}A_{k+1}\geq\tfrac{2}{3},\quad\tfrac{1}{L+M}\geq\tfrac{4}{C(\sqrt{2}-1)^{2}}.$
Therefore, we have
$\tfrac{\left(\sqrt{2}-1\right)A_{k+1}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}{\geq}\tfrac{4a_{k}^{\frac{3}{2}}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{3\sqrt{C}}$
(56) $\geq-\min\limits_{x\in\mathbb{R}^{n}}\left\{a_{k}\langle\nabla f(x_{k+1}),x-\nu_{k}\rangle+\tfrac{C\|x-\nu_{k}\|^{3}}{12}\right\}.$
Substituting (55) and (A.1) into (A.1), $\mathcal{R}_{k+1}^{1}$ holds. Using $\mathcal{R}_{k}^{1}$ and $\mathcal{R}_{k}^{2},$ we have
$A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}A_{j}\tfrac{\|\nabla f(x_{j})\|^{\frac{3}{2}}}{\sqrt{L+M}}\overset{\text{$\mathcal{R}_{k}^{1}$}}{\leq}f_{k}^{*}\leq f_{k}(x)\overset{\text{$\mathcal{R}_{k}^{2}$}}{}$
${\leq}A_{k}f(x)+\left[\left(\tfrac{1}{2\sqrt{2}}+\tfrac{1}{6}\right)(L+M)+\tfrac{C}{6}\right]\|x_{0}-x\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}.$
Substituting the choice of $A_{k},C,M,$ and $x=x^{*}$ concludes the proof.
For a general $p$-th order accelerated tensor method, we can utilize [25, Corollary 1] and analyze a modified accelerated tensor method to derive a slow gradient complexity $\mathcal{O}(1/\varepsilon^{p})$, which satisfies 1 and thus can be used as a subroutine in the high-order AR framework. The proof follows the same structure as in [25] by modifying the estimating sequence and is analogous to the accelerated CNM in how the gradient-norm output is obtained; therefore, we omit the details for simplicity.